OpenStudy (anonymous):

how do I solve w - - 8=-3 6

OpenStudy (zepp):

Send 8 to the right side by adding 8. \(\large \frac{w}{6} = -3+8\) \(\large \frac{w}{6} = 5\)

OpenStudy (zepp):

Then to find w, you simply get rid of 6 by dividing by 6 to both side.